OpenStudy (anonymous):

a rectangle has an area of 21 and a perimeter of 20. what is the length of its diagonal?

OpenStudy (akashdeepdeb):

This is a fun Question!! :D Okay so if area = 21 Then l * b = 21 peremeter = 20 = 2*(l+b) Or (l+b) = 10 So we got 2 equations.. (l+b) = 10 lb = 21 HINT: Use (l+b)^2 = l^2 + b^2 + 2lb Diagonal is found by pythagorus theorem!! :D